If sales are $820,000, variable costs are 55% of sales, and operating income is $260,000, what is the contribution margin ratio?

________2. A firm operated at 90% of capacity for the past year, during which fixed costs were $420,000, variable costs were 40% of sales, and sales were $1,000,000. Operating profit was:

________3. Bryce Co. sales are $914,000, variable costs are $498,130, and operating income is $196,000. What is the contribution margin ratio?

________4. A firm operated at 80% of capacity for the past year during which fixed costs were $330,000, variable costs were 70% of sales, and sales were $1,000,000. Operating profit (loss) was:

________5. Zeke Company sells 25,000 units at $21 per unit. Variable costs are $10 per unit, and the fixed costs are $75,000. The contribution margin ratio and the unit contribution margin are:

________6. If the fixed costs are $1,200,000, the unit selling price is $240, and the unit variable costs are $110, what is the amount of sales required to realize an operating income of $200,000?

________7. If the fixed costs are $500,000, the unit selling price is $55, and the unit variable costs are $30, what is the break-even sales (units) if the fixed costs are increased by $80,000?

________8. Reynold's Grocery has fixed costs of $350,000, the unit selling price is $29, and the unit variable costs are $20. What is the break-even sales (units) if the variable costs are decreased by $4?

________9. Johnson's Plumbing's fixed costs are $700,000 and the unit contribution margin is $17. What amount of units must be sold in order to realize an operating income of $100,000?

________10. If the fixed costs are $561,000 and the unit contribution margin is $8.00, what is the break-even point in units if the variable costs are decreased by $0.50 a unit?
Answer #1

1. Contribution margin ratio = 100% - Variable cost ratio

= 100% - 55%

= 45%

Option A is the answer

2. Operating profit = Contribution margin - Fixed cost

= (1,000,000*60%) - 420,000

= 180,000

Option A is the answer

3. Contribution margin ratio = (Sales - Variable cost) /Sales

= (914,000-498,130)/914,000

= 45.5%

Option D is the answer

4. Operating profit = Contribution margin - Fixed cost

= (1,000,000*30%) - 330,000

= (30,000)

Option B is the answer